# General Flow as Foundation Affordance for Scalable Robot Learning

### v1

@@ Abstract
We address the challenge of acquiring real-world manipulation skills with a scalable framework. Inspired by the success of large-scale auto-regressive prediction in Large Language Models (LLMs), we hold the belief that identifying an appropriate prediction target capable of leveraging large-scale datasets is crucial for achieving efficient and universal learning. Therefore, we propose to utilize flow, which represents the future trajectories of 3D points on objects of interest, as an ideal prediction target in robot learning. To exploit scalable data resources, we turn our attention to cross-embodiment datasets. We develop, for the first time, a language-conditioned prediction model directly from large-scale RGBD human video datasets. Our predicted flow offers actionable geometric and physics guidance, thus facilitating stable zero-shot skill transfer in real-world scenarios. We deploy our method with a policy based on closed-loop flow prediction. Remarkably, without any additional training, our method achieves an impressive 81% success rate in human-to-robot skill transfer, covering 18 tasks in 6 scenes. Our framework features the following benefits: (1) scalability: leveraging cross-embodiment data resources; (2) universality: multiple object categories, including rigid, articulated, and soft bodies; (3) stable skill transfer: providing actionable guidance with a small inference domain-gap. These lead to a new pathway towards scalable general robot learning. Data, code, and model weights will be made publicly available.
I Introduction
@@
We aim to reveal a potential pathway for replicating the success of Large Language Models (LLMs) in the domain of robot learning. Specifically, we are interested in developing a new framework that enables scalable learning for robot manipulation. With more data and larger model training in the future, this framework has the potential to progressively enhance the capabilities of robots, i.e., the scaling law that has been observed in LLMs [82]. Inspired by the LLMs training paradigm [14], we believe that two key elements contribute to their strong generalization abilities: (1) a vast training dataset with a small inference domain gap, such as all texts from the internet in LLMs, and (2) a foundational prediction task with appropriate supervision signals, such as text-token prediction in LLMs. How can we translate these elements into robot learning?
@@
Confronted with the challenges of collecting real-world robot data [46, 67], we pivot towards large-scale human datasets. These data resources guarantee scalability and a small inference domain-gap (no simulation-to-reality problem), key ingredients for effective generalization. Moreover, human manipulation data provides a vast, real-world dataset rich in diverse physics interactions and dynamic behaviors that closely align with robot manipulation. The next step is to identify a foundational prediction target conducive to widespread downstream tasks. We propose affordance for this role. Rooted in Gibson’s theory [31], affordance concentrates on the potential actions associated with an object, remaining neutral to specific manipulators. This characteristic positions affordance as a cornerstone in cross-embodiment robot learning.
@@
What affordance format will lead to a foundation prediction target that is universal for object categories and provides actionable geometric and physics guidance for downstream applications? In this paper, we propose General Flow as Foundation Affordance (as shown in Figure 1) to achieve this goal. This affordance elucidates the future trajectories of 3D points on the object of interest. Our key observation is that predicting keypoint motion is an efficient way to express geometric and physics information. Take the task of ‘open Safe’ as an instance (in the middle of Figure 1): the general flow represents future positions of points on the safe. Then, a robot can infer the safe’s articulation by noting a static flow on the body and a moving flow on the door. It can also gain a resilient motion primitive for the opening skill by following the door’s flow. We term our affordance “general flow” due to its capability for universal robot learning: (1) scalability: leveraging different embodiment data resources, e.g., humans and different robots; (2) universality: multiple object categories, including rigid, articulated, and soft bodies, with a large number of potential applications (see Section II-B). (3) stable skill transfer: providing actionable geometric and physics information with a small inference domain-gap, even sufficient for zero-shot execution.
@@
In this paper, a novel framework is proposed to leverage general flow as the training target for foundational affordance learning. We first develop pipelines to extract 3D flow labels directly from RGBD human video datasets. We find prediction of dense flow in real-world scene point clouds remains a formidable challenge, primarily due to the variability of trajectory scales and the need to enhance robustness in zero-shot scenarios. To address these issues, we employ scale-aware strategies in both the data and model aspects, complemented by augmentation techniques that focus on embodiment occlusion (human hand and robot arm) and query point sampling (3D points on objects of interest), thereby boosting zero-shot stability. Remarkably, our model (named “ScaleFlow”), with fewer parameters, surpasses existing methods, setting a strong baseline for future research. Moreover, our system, trained at scale, demonstrates notable competencies such as language-driven semantic control, resilience to label noise, and spatial commonsense understanding.
@@
To showcase the full potential of general flow affordance, we elect to tackle one of its most challenging applications: zero-shot cross-embodiment execution. Implementing a straightforward heuristic policy derived from closed-loop flow prediction, we evaluate our approach on a Franka-Emika robot in a real-world setup. Distinct from prior methodologies [4], without any additional training, our system accomplishes stable zero-shot human-to-robot skill transfer. Despite the simplicity of the derived policy, our approach, fueled by the rich actionable geometric and physics guidance of general flow affordance, notches an impressive 81% average success rate in 18 diverse tasks across 6 scenes, covering multiple categories of object types like rigid, articulated, and soft bodies. It is also noteworthy that our approach capably handles complex environmental challenges, such as robot segmentation errors, novel categories, and various robot configurations (initial gripper positions, grasp states, and manipulation directions), to a considerable extent. These findings highlight the transformative potential of general flow in spearheading scalable general robot learning.
In summary, our contributions can be concluded as follows:
We introduce the framework of General Flow as Foundation Affordance, substantiating its feasibility and effectiveness, which is a new pathway towards scalable robot learning.
We propose a robust, scale-aware algorithm that utilizes 3D flow labels derived from RGBD human video datasets, achieving remarkable accuracy in predicting complex real-world flow scenarios.
We apply a simple heuristic policy, based on our model, to a Franka-Emika robot, successfully enabling stable zero-shot human-to-robot skill transfer across various object categories. This results in an impressive 81% success rate in 18 tasks across 6 scenes, marking a significant milestone in the real-world application of flow-based methods.
II Related Work
II-A Universal Robot Learning for General Manipulation
Real-World General Robot Learning Research in general-purpose robotic manipulation in real-world settings is constantly evolving, with a focus on integrating Large Language Models (LLMs) for high-level planning [13, 44, 37, 43, 19, 41] and exploring direct actionable guidance through LLMs [45], albeit with challenges due to overlooked physics dynamics. The development of large models for direct low-level control [11, 12, 9, 66] faces scalability issues due to intensive data requirements [46, 67]. This highlights the need for a training framework that achieves a balance between actionable output and scalability. In this work, we achieve this through a universal robot learning approach based on general flow prediction.
Embodiment-Agnostic Framework for Robot Learning To leverage large-scale, cross-embodiment data resources [62, 28, 16, 55, 27], multiple embodiment-agnostic frameworks [32] are proposed for robot learning. Prior works [65, 56, 57, 90] employ large-scale visual pre-training to develop embodied-aware pretrained representations, but these demonstrate limited generalization [58, 38, 42]. Alternative approaches seek to derive action signals from image or video generation [20, 35, 25, 23, 21, 50, 7], but these methods are resource-intensive and often yield redundant information. Affordances extracted from simulators [88, 63, 81, 30, 89, 29] are another focus, yet they struggle with the significant sim-to-real domain-gap, particularly in 3D environments. Recent efforts [3, 4, 61] attempt to directly acquire geometric-aware structured information in human video but require burdensome in-lab training. Bharadhwaj et al. [8] use hand poses as a bridge for guiding robot manipulation, but the limited geometric information extraction leads to unstable performance. Instead, we leverage 3D flow-based affordance to achieve reliable solutions. This leads to a stable skill transfer with an impressive 81% success rate in the real world.
Keypoints and Flow for Robot Learning Systems Utilizing keypoints and flow efficiently reveals the fine-grained geometric properties of the physical world, prompting researchers to harness their power in robot learning systems. Previous studies [59, 26, 24, 78, 91, 79, 60, 68] use keypoints as state descriptors for robot learning. The potential of flow prediction has also been noted [86, 76, 17, 2, 80]. However, these approaches either depend on embodiment-specific data, limiting their scalability [69, 74], or are simulation-based, facing significant sim-to-real domain-gap due to imperfect real-world RGBD point cloud generation [22, 93]. In this paper, we extend the work of Seita et al. [74], Eisner et al. [22] into a more general version (termed “general flow”) both in data resources and downstream applications. We acquire a large-scale 3D flow prediction model directly from RGBD human video datasets and achieve stable zero-shot skill transfer. Besides, we propose a novel scale-aware architecture design and robust augmentation techniques to address the challenges of real-world flow prediction. Compared with previous models [70, 71, 72, 22, 74, 69], our methods achieve a significant performance enhancement.
II-B Potential Application for General Flow Prediction
General flow prediction (formally defined in the next section) can provide a foundational affordance prior [31, 15] for robot manipulation. Here, we outline some representative potential downstream applications. In perception, using algorithms like clustering, general flow can facilitate coarse part-level segmentation of objects [64] and serve as a prior for pose prediction and tracking [95, 83, 84]. For planning, robot motion planning can be executed [59, 26, 97, 22] based on flow prediction. Additionally, general flow can act as a strong policy prior [92] for robot learning, applicable in both imitation [74, 85] and reinforcement learning [4, 3]. In safety applications, the robot can detect anomalies by tracking keypoints and comparing them with predicted flow [36]. Finally, general flow can also provide semantic priors for other embodied tasks, such as Human-Object Interaction (HOI) synthesis [52].
Verifying all these applications is beyond the scope of this paper. Instead, we focus on the most challenging task: stable real-world cross-embodiment zero-shot skill transfer, a task not yet achieved by any existing flow-based work.
III General Flow as Foundation Affordance
III-A General Flow Affordance
Manipulation tasks typically consist of functional grasp and subsequent motion [1, 47]. In this paper, we mainly focus on the affordance [31] of later. We introduce “general flow” as an affordance that provides comprehensive, actionable guidance in terms of geometry and physics for downstream manipulation tasks:
@@ III-A General Flow Affordance
Definition of General Flow: Given a perception observation (from any embodiment) and a task instruction , for 3D query points in space, the general flow represents the trajectories of these points over future timestamps.
In this work, we use point clouds from real-world RGBD camera streams as our perception state , eliminating sim-to-real transfer domain-gap concerns. We term our affordance “general flow” to emphasize its broad applicability across different embodiments, object categories, and downstream applications. Next, we delve into its distinct properties to highlight its potential and effectiveness.
III-B Properties of General Flow
We explore the advantageous properties of the general flow affordance (see Figure 2) to demonstrate its capability as a bridge to a scalable, general robot learning framework. It provides several key benefits:
Scalability: General flow allows for direct utilization of cross-embodiment data, such as human video datasets, circumventing the challenge of accessing large volumes of real robot data [74, 66]. Moreover, representing physics as future trajectories is a resource-efficient abstraction of motion dynamics, especially compared with full video generation [20, 50].
Universality: General flow represents a unified abstraction of physical dynamics across multiple object categories, e.g., rigid, articulated, and soft bodies [29]. It also provides support for a wide range of applications (Section II-B). Additionally, its predictions are contingent on language instructions, enabling the execution of various behaviors within a single scene.
Stable Skill Transfer: This benefit arises from two aspects. First, general flow offers richer geometric and physical guidance, especially compared to pretrained representations [65, 56] and coarse motion trends [4, 8]. Second, its reliance on real-world data eliminates any sim-to-real domain-gap issue. [22, 88].
Considering these points, we posit that general flow offers a scalable prediction target for foundation robot learning, similar to “text token” in Large Language Models (LLMs). Given that scaling up has led to strong generalization and emergent phenomena in LLMs [82, 14], we expect similar progressive enhancements in robot capabilities through larger-scale training. In the future, we aim to achieve this by harnessing larger RGBD datasets [34] recently released or by combining RGB video resources [16, 33] with depth estimation techniques [10, 50].
IV Embodiment-Agnostic and Scale-Aware General Flow Prediction
In this section, we propose a framework for general flow prediction that is agnostic to specific embodiments, which is outlined in Figure 3. We first design pipelines to extract flow labels from RGBD human video datasets. To manage the variable scales of trajectories and account for real-world noise, we integrate key designs that enhance the model’s scale-awareness and robustness in predictions. For more details on the label pipeline and training insights, please refer to Appendix -B,-C,-D.
IV-A General Flow Label Acquisition
We introduce methods for acquiring general flow labels from two types of cross-embodiment datasets. All tools and pipelines will be open-source to benefit future research.
From 3D Annotated Datasets [51, 28, 55]: Utilizing the detailed 3D labels from these datasets, we first randomly sample points within the active object and then calculate its future position using ground-truth pose and camera parameters.
From Annotation-Free RGBD Videos [27, 34]: We first perform Human-Object-Interaction (HOI) segmentation [75, 96] to obtain the active object mask. Points are then sampled within this mask, and the future 2D trajectory is tracked using the Tracking Any Point (TAP) tools [48]. The 3D label of the general flow is determined through back projection in both the spatial and temporal dimensions.
To reduce the effect of noise in the annotations and the pipeline, multiple techniques and filters are employed. Additionally, we retain the hand mask for potential use in subsequent training augmentations.
IV-B Scale-Aware Prediction Model
@@
Our model processes natural language instructions , scene point cloud features (comprising points with XYZ+RGB attributes), and spatial query points (comprising points with XYZ attributes). The aim is to predict a trajectory set, or “flow”, denoted as . For the -th query point , its trajectory is defined as , with the absolute position at time represented as for . Initially, is set to the input position of the query point . We observe enhanced performance when predicting relative displacements rather than absolute positions. Thus, our refined goal is to predict for . The trajectory length for each query point is defined as .
A primary challenge in real-world flow prediction is the significant variance in trajectory lengths across different query points. For instance, in the “open Safe” task, the trajectories of points on the door are substantially longer than those on the safe body. To address this, we apply Total Length Normalization (TLN) to uniformly rescale trajectories. For the original prediction target , we define the scale and normalized target as:
Our ablation study demonstrates that TLN yields the best performance compared with other normalization methods (Appendix -F), leading to its adoption in subsequent experiments. The original prediction target can be easily reconstructed from the predicted values of and .
Next, we describe our model’s architecture (Figure 4). To facilitate multimodal control, we integrate instruction semantics early in the process. Instructions are converted into semantic features using a CLIP [73] encoder, then their dimensions are reduced via MLPs (to ) to align with point features. Features of scene point clouds include 3D positions and RGB values, while query point clouds substitute RGB values with a learnable embedding (it serves as a “query identifier” and remains the same for all query points). We first concatenate aligned text features with point cloud features, then concatenate the features of scene points and query points, forming merged point cloud features . The merged features are processed through a PointNeXt [72] backbone with a segmentation head to extract geometric features. The features of query points serve as condition variables for a conditional VAE [49, 40], generating the final predictions and . For consistent scene semantics, a single latent variable is sampled for all query points within the same scene during inference.
IV-C Loss Calculation with Scale Rebalance
The final loss function comprises trajectory reconstruction loss , scale regression loss , and VAE KL-divergence loss . To minimize cumulative error, we also incorporate an MSE loss for the recovered accumulative shift. Thus, the total loss is expressed as:
Our experiments underscore the importance of adequately weighting by setting and , which is crucial for optimal performance.
IV-D Augmentations for Zero-Shot Robustness
In light of the complex environmental challenges encountered in zero-shot real-world deployments, we propose two technical augmentations to boost zero-shot generalization robustness:
Hand Mask (HM) Augmentation: We encounter occlusions from human hands in our training data while facing occlusions from robot arms during deployments. Therefore, it is crucial to enhance the model’s resilience to embodiment occlusions. To achieve this, we manipulate the presence of points on the hand in the input scene point clouds. We choose one of three rules, with probabilities =0.5, =0.2, and =0.3: (1) deleting all hand points; (2) keeping all hand points; and (3) sampling a random anchor point on the hand and retaining only points with a distance from the anchor greater than 12cm.
Query Points Sampling (QPS) Augmentation: Different downstream applications may require varying query point sampling methods. Consequently, our model must be adaptable to various query point distributions. We achieve this by augmenting the training process. In each training iteration, we select a subset of available query points using one of two rules, based on probabilities =0.7, =0.3: (1) complete random sampling; (2) randomly selecting an anchor query point and then choosing a specific number of points closest to the anchor.
V General Flow Prediction Experiment
V-A Experimental Setting
@@
Dataset: For rigid and articulated objects, we utilize the HOI4D dataset [55] to train our general flow prediction model. This extensive RGBD video dataset includes 16 categories and 800 objects, encompassing 44.4 hours of recording. It provides comprehensive 3D labels, such as active object segmentation, 3D pose, and camera parameters. To further explore general flow in soft object manipulation, we collect RGBD videos for the “fold clothes” task using the RealSense D455 camera, comprising 6 garments, 30 rollouts, and 605 extracted clips.
| Model | ADE | FDE | ADE-H | FDE-H | Param(M) |
| ResNet18 | 0.0754 | 0.1071 | / | / | 13.160 |
| R3M(frozen) | 0.0755 | 0.1056 | / | / | 11.861 |
| R3M(finetune) | 0.0754 | 0.1069 | / | / | 11.861 |
| VAT-MART | 0.0716 | 0.1220 | 0.0717 | 0.1220 | 1.577 |
| VIT-B-224 | 0.0681 | 0.0948 | / | / | 86.614 |
| PointNeXt-B | 0.0396 | 0.0537 | 0.0392 | 0.0529 | 4.134 |
| PointNeXt-L | 0.0383 | 0.0516 | 0.0380 | 0.0512 | 15.583 |
| ScaleFlow-S | 0.0374 | 0.0501 | 0.0372 | 0.0498 | 0.906 |
[…]
| ScaleFlow-L | 0.0355 | 0.0470 | 0.0352 | 0.0467 | 17.088 |
Baseline: Given the absence of an identical problem setting in previous work, we adapt three types of relevant work to our setting:
We refer to our model as “ScaleFlow” in subsequent discussions. For 3D backbones and our model, we train multiple versions with varying model sizes. More details are available in Table I and Appendix -C. To ensure a fair comparison, scale rebalance, HM augmentation and QPS augmentation are applied to all baselines.
Training Details: We utilize 1.5s video clips as our training data and set the time steps of general flow to 3 for all data sources. The dataset is divided into training, validation, and test sets in an 80%, 10%, 10% ratio, resulting in 51693, 6950, and 6835 clips respectively, with no identical object instances across sets. Each sample consists of 2048 scene points sampled in an 808080 cube space around the center of the flow start points using the furthest point sampling (FPS) algorithm. During training, we randomly sample 128 query points, while for validation, 512 points are randomly sampled. Given that our datasets include ground-truth labels for object parts, we distribute 512 query points across each part equally during testing to enhance evaluation effectiveness. It’s important to note that we do not use any part labels in model validation and real-word testing.
We train all the models with a batch size of 128, utilizing an AdamW optimizer and a learning rate of 0.001. For validation and testing, we set the batch sizes to 256. All models are trained for 200 epochs. Notably, the base version of ScaleFlow (ScaleFlow-B), which has 5.62M parameters, completes its training in 10 hours on a single NVIDIA GeForce RTX 3090 GPU.
Evaluation Metrics: We use 3D Average Displacement Error (ADE) and Final Displacement Error (FDE) in meters [5, 53] as evaluation metrics. For VAE-dependent models, metrics are averaged over 10 samplings. We also test robustness to hand occupancy in inputs across all 3D models, which are denoted as ADE-H and FDE-H.
V-B Prediction Results
The results in Table I demonstrate that our model achieves superior performance on all metrics, even with fewer parameters. Overall, 3D models are superior to 2D models, indicating the importance of 3D geometry. Notably, with appropriate augmentations, the model is robust for hand occupancy. We show visualization of flow prediction in Appendix -E. Furthermore, our ablation study (Appendix -F) results quantitatively illustrate the effectiveness of all designs.
Through large-scale training, our model not only captures rich semantic information but also becomes adeptly controllable through language modality. As depicted in Figure 5(a)(b), our model demonstrates the capability to predict varied flows for identical input point clouds when provided with different instructions. Furthermore, it is remarkably robust to label noise. Figure 5(c)(d) showcases two instances of this resilience: despite severe label noise (notable deviation in “open Safe” and near-static in “pickup Toy Car”), our model accurately predicts the correct trend. Additionally, our model gains spatial commonsense through scalable training. It dynamically adjusts its prediction scale in response to the spatial relationships of objects, such as ending on the table and scaling up for longer distances, as seen in Figure 5(e)(f). All these emerging phenomena reveal the benefits of large-scale training.
VI Zero-shot Real World Manipulation

### v2

@@ Abstract
We address the challenge of acquiring real-world manipulation skills with a scalable framework. We hold the belief that identifying an appropriate prediction target capable of leveraging large-scale datasets is crucial for achieving efficient and universal learning. Therefore, we propose to utilize 3D flow, which represents the future trajectories of 3D points on objects of interest, as an ideal prediction target. To exploit scalable data resources, we turn our attention to human videos. We develop, for the first time, a language-conditioned 3D flow prediction model directly from large-scale RGBD human video datasets. Our predicted flow offers actionable guidance, thus facilitating zero-shot skill transfer in real-world scenarios. We deploy our method with a policy based on closed-loop flow prediction. Remarkably, without any in-domain finetuning, our method achieves an impressive 81% success rate in zero-shot human-to-robot skill transfer, covering 18 tasks in 6 scenes. Our framework features the following benefits: (1) scalability: leveraging cross-embodiment data resources; (2) wide application: multiple object categories, including rigid, articulated, and soft bodies; (3) stable skill transfer: providing actionable guidance with a small inference domain-gap. Code, data, and supplementary materials are available https://general-flow.github.io/.
Keywords: Flow, Transferable Affordance, Human Video
1 Introduction
@@
We aim to develop a new framework that enables scalable learning for robot manipulation. With more data and larger model training in the future, this framework has the potential to progressively enhance the capabilities of robots, i.e., the scaling law that has been observed in LLMs [1]. Inspired by the LLMs training paradigm [2], we believe that two key elements contribute to their strong generalization abilities: (1) a vast training dataset with a small inference domain gap, such as all texts from the internet in LLMs, and (2) a foundational prediction task, such as text-token prediction in LLMs. How can we translate these elements into robot learning?
@@
Confronted with the challenges of collecting real-world robot data [3, 4], we pivot towards large-scale human video datasets. These data resources guarantee scalability and a small inference domain-gap (no simulation-to-reality problem), key ingredients for effective generalization. Moreover, human manipulation data provides a vast, real-world resource rich in diverse physics interactions and dynamic behaviors that closely align with robot manipulation. The next step is to identify a proper prediction target. We propose affordance for this role. Rooted in Gibson’s theory [5], affordance concentrates on the actions associated with an object, remaining neutral to manipulators. This characteristic positions affordance as a cornerstone in humtn-to-robot skill transfer.
@@
What affordance format will lead to a foundation prediction target that is universal for object categories and provides actionable guidance for robot manipulation? In this paper, we propose General Flow as Foundation Affordance (as shown in Figure 1) to achieve this goal. This affordance elucidates the future trajectories of 3D points on the object of interest. Take the task of ‘open Safe’ as an instance (in the middle of Figure 1): the general flow represents future positions of points on the safe. Then, a robot can gain a resilient motion primitive for the opening skill by following the door’s flow.
Previous works [6, 7, 8, 9] have attempted to extract 3D flow representations from either simulation or real robot data. However, these approaches suffer from domain transfer gaps [6] or have limited scalability [7]. In contrast, general flow leverages real-world, scalable data resources, i.e., human videos, thereby eliminating both the sim-to-real domain gap and the dependence on burdensome robot data collection. The diversity of backgrounds, objects, and human behaviors in these videos also significantly enhances the robustness of real robot execution. We term our affordance “general flow” due to its capability for universal robot learning: (1) scalability: leveraging scalable human data resources; (2) wide application: multiple object categories, including rigid, articulated, and soft bodies. (3) stable skill transfer: providing actionable guidance with a small inference domain-gap, even sufficient for zero-shot execution.
@@
In this paper, we first develop pipelines to extract 3D flow labels directly from RGBD human video datasets for model training. We find prediction of dense flow in real-world scene point clouds remains a challenge, primarily due to the variability of trajectory scales and the need to enhance robustness in zero-shot scenarios. To address these issues, we employ scale-aware strategies in the model aspect, complemented by augmentation techniques that focus on embodiment occlusion (human hand and robot arm) and query point sampling (3D points on objects of interest), thereby boosting zero-shot stability.
@@
Implementing a straightforward heuristic policy derived from closed-loop flow prediction, we evaluate our approach on a Franka-Emika robot in a real-world setup. Without any in-domain finetuning, our system accomplishes stable zero-shot human-to-robot skill transfer. Fueled by the rich actionable guidance of general flow affordance, with only one model, our system notches an impressive 81% average success rate in 18 diverse tasks across 6 scenes, covering multiple categories of object types like rigid, articulated, and soft bodies. These findings highlight the transformative potential of general flow in spearheading scalable general robot learning.
2 Problem Formulation
The affordance [5] typically consist of functional grasp and subsequent motion [10, 11]. While robust functional grasping has been extensively explored [12, 13, 14], a general method for providing prior guidance for post-grasp movements remains a challenge, which is the primary focus of our paper.
Understanding the semantics and geometry of post-grasp motion for accurate predictions is complex. For example, in the ”open Safe” task, the model must infer that the door’s point trajectory rotates around an implicit axis, while points on the base remain static. Additionally, it must note that the trajectories of closer points to the axis require a smaller predicted scale or length. To handle these, we introduce “general flow” as an affordance that provides actionable guidance for downstream manipulation:
@@ 2 Problem Formulation
Definition of General Flow Given a perception observation (from any embodiment) and a task instruction , for 3D query points in space, the general flow represents the trajectories of these points over future timestamps.
@@
Detailed Formulation In this work, we use point clouds from real-world RGBD camera streams as our perception state . Our model processes natural language instructions , scene point cloud features (comprising points with XYZ + RGB attributes), and spatial query points (comprising points with XYZ attributes). The aim is to predict a trajectory set, or “flow”, denoted as . For the -th query point , its trajectory is defined as , with the absolute position at time represented as for . Initially, is set to the input position of the query point .
We term our affordance “general flow” to emphasize its broad applicability across different embodiments (from any embodiment to any robot embodiment, e.g. from human to Franka-Emika) and object categories (e.g. rigid, articulated and soft bodies).
3 Embodiment-Agnostic Flow Prediction
We propose a framework for general flow prediction that is agnostic to specific embodiments, which is outlined in Figure 2. We first design pipelines to extract flow labels from RGBD human video datasets. To manage various scales of trajectories and account for real-world noise, we integrate key designs that enhance the model’s scale-awareness and robustness in predictions. For the first time, we develop a language-conditioned 3D flow prediction model directly from large-scale RGBD human video datasets.
3.1 General Flow Label Acquisition
@@
We extract general flow labels from two types of cross-embodiment datasets. For rigid and articulated objects, we utilize the HOI4D dataset [15] to train our general flow prediction model. This extensive RGBD video dataset includes 16 categories and 800 objects, encompassing 44.4 hours of recording. It provides comprehensive 3D labels, such as active object segmentation, 3D pose, and camera parameters. To further explore general flow in soft object manipulation, we collect RGBD videos for the “fold clothes” task using the RealSense D455 camera, comprising 6 types of clothes, 30 trajectories, and 605 extracted clips.
From HOI4D Datasets[15]: Utilizing the detailed 3D labels from these datasets, we first randomly sample points within the active object and then calculate its future position using ground-truth pose and camera parameters.
From Collected RGBD Videos: We first perform Human-Object-Interaction (HOI) segmentation [16, 17] to obtain the active object mask. Points are then sampled within the mask, and the future 2D trajectory is tracked using the Tracking Any Point (TAP) models [18]. The 3D label of the general flow is computed through back projection in both the spatial and temporal dimensions.
To reduce the effect of noise in the annotations and the pipeline, multiple filtering techniques are employed. Additionally, we retain the hand mask for potential use in training augmentations. More details are shown in Appendix B.
| Model | ADE | FDE | PM |
| ResNet18 | 7.54 | 10.71 | 13.2 |
| R3M(frozen) | 7.55 | 10.56 | 11.9 |
| R3M(finetune) | 7.54 | 10.69 | 11.9 |
| VAT-MART | 7.16 | 12.20 | 1.6 |
| VIT-B-224 | 6.81 | 9.48 | 86.6 |
| PointNeXt-B | 3.96 | 5.37 | 4.1 |
| PointNeXt-L | 3.83 | 5.16 | 15.6 |
| ScaleFlow-S | 3.74 | 5.01 | 0.9 |
| ScaleFlow-B | 3.58 | 4.77 | 5.6 |
| ScaleFlow-L | 3.55 | 4.70 | 17.1 |
3.2 Scale-Aware Flow Prediction
We observe enhanced performance when predicting relative displacements rather than absolute positions. I.e., we predict instead of . The illustration of the model architecture is shown in Figure 3. Following the description in Section 2, we first use CLIP [19] encoder to convert instructions into semantic features. Then their dimensions are reduced via MLPs (to dimensions) to align with point features. We first concatenate aligned text features with point cloud RGB+XYZ features, and then concatenate the features of scene points and query points, forming merged point cloud features . The merged features are processed through a PointNeXt [20] backbone with a segmentation head to extract geometric features. Since humans may perform different action trajectories for the same task and scene, we utilize a VAE [21, 22] to capture this multimodal property. The features of query points serve as condition variables for a conditional VAE, generating the final predictions. We leave more details of architecture in Appendix C.1.
A primary challenge in real-world flow prediction is the significant variance in trajectory lengths across different query points within a task. For instance, in the “open Safe” task, the trajectories of points on the door are substantially longer than those on the safe body. To address this, we apply the Total Length Normalization (TLN) to uniformly rescale trajectories. We define trajectory scale for each query point as . For the original prediction target , we define the normalized target as:
Then the VAE predicts the scale and normalized trajectory separately. Our ablation study demonstrates that TLN yields the best performance compared with other normalization methods (Appendix F.1).
3.3 Training Flow Prediction Model
The final loss function comprises trajectory prediction loss , scale regression loss , and KL-divergence loss of the VAE[21]. To minimize cumulative error, we also incorporate an MSE loss for the recovered accumulative shift. Thus, the total loss is expressed as:
In light of the complex environmental challenges encountered in zero-shot real-world deployments in later sections, we propose two technical augmentations to boost zero-shot generalization robustness. (1) Hand Mask Augmentation (HMA): to enhance resilience to embodiment occlusions, we drop out the hand points in the input scene point clouds with some probabilities. (2) Query Points Sampling (QPS): to adapt to varying query point distributions required by different applications, we augment training by switching query point distributions. For more details of the augmentations and training procedures, please refer to Appendix C.2, C.3. Additionally, an ablation study is presented to verify the effectiveness of all design elements in Appendix F.
3.4 Evaluation
The collected general flow dataset is divided into training, validation, and test sets in an 80%, 10%, 10% ratio, with no identical object instances across sets. We adopt three types of models as our baselines (demonstrated in Appendix D). (1) 2D model: ResNet[23], ViT[24], R3M[25]. (2) Point clouds affordance model: VAT-MART[10]. (3) 3D model: PointNeXt[20], which is an improved version of the PointNet++[26] used in Flowbot3D[6] and ToolFlowNet[7]. We refer to our model as “ScaleFlow” in subsequent discussions. We trained multiple versions of the PointNeXt-based models, each with different PointNeXt backbone sizes.
We use 3D Average Displacement Error (ADE) and Final Displacement Error (FDE) in centimeters [27, 28] as evaluation metrics. For VAE-dependent models, metrics are averaged over 10 samplings. The results in Table 1 demonstrate that our model achieves superior performance on all metrics, even with fewer parameters. We show the visualization of flow prediction in Figure 4. Moreover, our system, trained at scale, demonstrates notable emergent properties such as robustness to hand occlusion, language-driven semantic control, resilience to label noise, and scale spatial adaption. We demonstrate these properties in Appendix E.1 and validate them in real-robot experiment in Appendix E.2.
4 Zero-shot Real World Manipulation